# Towards a Formal Foundation for Blockchain ZK Rollups

### v2

@@ 3.1 Threat Model
Users. Users interact with the L2 either by submitting their transactions directly to the L2 network, i.e., the rollup operators, or by submitting them through the rollup’s contract on the L1 via the forced queue.
@@
Rollup administrators. In addition to rollup operators, some entities have administrator capabilities for the L1 blockchain smart contracts of the L2. Administrators can be white-listed or elected via governance mechanisms. These administrators have the ability to request system upgrades or update blacklists. In practice, currently, the administrator capabilities are assigned to a multi-sig that could be controlled by some distinct entities. The mechanisms we describe in this paper should maintain security even if administrators are malicious.
@@
  • The cryptographic primitives (ZKPs) and their implementation are assumed to be secure [18], ensuring that no verifier accepts an invalid proof of execution. This assumption also extends to the correctness of any trusted setup procedures, the deployed verification keys on the L1 verifier contract, and the smart contract code responsible for proof verification.
  • Smart contracts operating L2 functionalities on the L1 blockchain are considered secure against conventional vulnerabilities, such as re-entrancy attacks [19], and are modifiable only through a well-defined upgrade process. In other words, in this paper, we focus on what may be called logical vulnerabilities.
@@
Strong Adversary Model. Any of the parties mentioned above — users, operators, and administrators — can become malicious. Users can perform DoS attacks, operators can engage in transaction censorship, and rollup administrators can execute malicious upgrades that could lead to the theft of all user assets. Despite these possibilities, the saving grace is that rollup contracts deployed on the underlying L1, together with the economic security that the L1 provides, allow us to use decentralized trust to address these issues. This comes down to the ability to analyze [20] these rollup contracts.
For example, zkSync Era has contracts such as Verifier.sol, which process validity proofs and can be upgraded by a multi-sig [^11] https://github.com/matter-labs/era-contracts. It is critical that this verifier can only be updated through a safe process because if it becomes malicious, users can lose their money. Hence, users should have a sufficient mechanism to exit the rollup in case of adversarial or questionable upgrades. While our model is extensible, this paper largely focuses on functional security, sidestepping DoS concerns.
3.2 User-centric Security Goals
@@
Weak liveness. A key goal is to ensure that any user transaction submitted to the forced queue is processed as long as the L2 continues to make progress. This guarantee is crucial to protecting users against potentially compromised or malicious L2s or even adding L1-like censorship resistance for centralized L2s. Moreover, our proposed design ensures that even when L2 contracts enforce transaction blacklisting, transactions cannot be halted if they have already been submitted to the forced queue. This mechanism is vital to maintain the integrity and safety of ZK-Rollups. In summary, weak liveness in the context of rollups means that either the L2 will eventually process user transactions submitted through L1 (i.e., inheriting the liveness of the L1), or the rollup will be frozen. In Section 4.3 we design forced queue mechanism which enforces weak liveness.
@@
4 Implementation in Alloy
In this section, we present the implementation details of our ZK Rollup model using the Alloy specification language. We first give a brief technical introduction to Alloy, followed by the Alloy model for a "strawman" ZK Rollup and three critical mechanisms, highlighting our reasoning and essential security properties. In Section 5, we demonstrate how to use our model to detect issues in flawed designs. Our approach directly encourages finding problems early, at the design stage, as opposed to later.
4.1 A Brief Introduction to Alloy
Alloy [15, 16] is a declarative modeling language that uses first-order relational logic to specify and check the properties of software systems. Data types in Alloy are represented as relations defined by signatures, which can be abstract or concrete. Abstract signatures categorize general types that are refined by more specific concrete signatures, creating a flexible and powerful type system. The language enforces the integrity of the model through facts, which are constraints that must always hold within the model. Functions and predicates in Alloy are named expressions and logical formulas, respectively; these allow the parametrization and reuse of complex logical constructs throughout the model.
@@ 4.1 A Brief Introduction to Alloy
The Alloy analysis is always limited to a specified scope. However, the small-scope hypothesis states that a high proportion of errors can be found by analyzing a system for all test inputs within some small scope [21].
This structured approach to modeling with Alloy ensures a rigorous verification of properties and behaviors in software systems, making it an invaluable tool in the development and analysis of reliable software architectures. In the upcoming sections, we present how we used Alloy to design and verify the correctness of critical elements of ZK-Rollup logic.
4.2 Strawman ZK Rollup Model
@@
    3 c.state.subseq[0,sub[#L1.finalized_state,1]] $=$ L1.finalized_state
    4 L1.commitments $=$ L1.commitments + c
@@
    4 c.state $=$ p.state
    5 c.diff $=$ p.diff
    6 c.state $=$ L1.finalized_state
@@
    8 L1.finalized_state $=$ L1.finalized_state.add[p.diff]
@@
7 c.diff $=$ p.diff and b $=$ c.diff
8 and c.state $=$ L1.finalized_state)
@@
Strictly speaking when we checked the property for 20 steps it still can fail for 21 steps. However, small-scope hypothesis states that a high proportion of errors can be found by checking a program for all test inputs within some small scope [21].
4.3 Forced Queue
@@
3 L1.forced_queue $=$ L1.forced_queue.add[f]
@@
5 and not L1.forced_queue.first.tx $=$ L1.forced_queue’.first.tx
@@
4.4 Forced Queue with Blacklisting
In recent years, transaction and address censorship has become a prominent issue for the Ethereum ecosystem. For example, 6% of Ethereum blocks are constructed to be OFAC-compliant [22], and Circle (USDC) maintains a list with blacklisted addresses [23]. Some L2 networks may choose to censor transactions to be OFAC compliant, potentially to be determined through their governance protocol. Next, we describe an “enshrined” blacklisting mechanism; this mechanism was meticulously designed, formalized, and checked with Alloy Analyzer to ensure it does not violate forced queue properties.
@@ 4.4 Forced Queue with Blacklisting
2 L1.forced_queue.first $=$ f
3 L1.blacklist $=$ L1.forced_queue.first.predicate
4 L1.forced_queue $=$ L1.forced_queue.delete[0]
@@
4.5 Forced Queue and Upgradeability
@@
  4. 4.
    Once the forced queue is cleared, the upgrade is deployed, and the rollup shifts to a new operational mode. During this timeout period, only forced queue events are processed, which pressures the L2 operator to swiftly handle the forced queue. Halting the acceptance of forced transactions risks the assurance of a permissionless exit via the forced queue mechanism. It is important to note that L2 users can still submit transactions through the L2 and receive pre-confirmations, but these cannot be finalized on the L1 until the forced queue is cleared.
@@
In summary, both blacklisting strategies have their trade-offs. More specifically, the blacklisting strategy through the forced queue (see Section 4.4) allows the ZK-Rollup to immediately prevent censored transactions to be added into the forced queue (however, it cannot forbid execution of the transactions which are already present in the queue). On the other hand blacklisting strategy through the upgradeability mechanism allows users to notice the upgrade and quit the ZK-Rollup within the predefined window of time without receiving a damage.
so the designer of the system can decide which mechanism they would like to implement in their roll-up.
@@
2 L1.ongoing_upgrade $=$ none
3 L1.ongoing_upgrade $=$ f
@@
3 not L1.blacklist $=$ L1.blacklist implies
@@
5 once (L1.ongoing_upgrade $=$ none and
6 L1.ongoing_upgrade $=$ x and
7 L1.blacklist $=$ is and
@@
2 (not L1.blacklist $=$ L1.blacklist’) implies
@@
Given all of the above properties, we arrive at the following guarantee:
@@
This mechanism ensures that users are not caught off-guard by upgrades and are guaranteed to be able to act upon upgrade (e.g., exit from L2 rollup) before the upgrade is enforced (if they react before the timeout).
Figure 3: Counter-example generated by Alloy, illustrating potential issues with the naive upgradeability mechanism. In this scenario, the blacklisting policy targets Input_1, leading to an L2 freeze because an input in the forced_queue cannot be processed. This occurs because the mechanism does not require processing all transactions from the forced queue prior to executing an upgrade. Note that despite an extended timeout period, adversarial users could still flood the forced queue with transactions, potentially leaving legitimate user transactions unprocessed before the upgrade occurs.
5 Design-time Testing with Alloy Models
Building on the implementation details and security properties described in the previous section, we explore how our Alloy models can be used for design-time testing of ZK-Rollups. We provide an infeasibility result and highlight a vulnerability in state-of-the-art upgradeability mechanisms.
@@ 5 Design-time Testing with Alloy Models
In this scenario for the rollup with upgradeability and blacklisting (Section 4.5), we start with an empty blacklist state and no queued commitments or proofs. We assert that, eventually, we would like to reach a state where there is an input x which is blacklisted but not finalized. After that we would like to eventually reach a state when the input x is finalized. This scenario is only possible if the rollup updates its blacklists two times where the first update blacklists x and the second update removes x from the blacklist and at the end x gets finalized. When we execute the run command, the Alloy Analyzer looks for models with up to seven instances of each signature which would satisfy the predicate. The Alloy Analyzer finds a satisfying instance that consists of a trace of length eight.
@@
Infeasibility of on-the-spot blacklisting. When processing a blacklisting policy update transaction (Section 2), transactions that precede it in the queue must be processed. This means that otherwise censored transactions would have to be processed. If blacklisting is applied immediately, it can generate a trace that gets the system stuck, breaking the guarantees of our system. To maintain the integrity of the forced queue, blacklisting should be enforced either through the forced queue (Section 4.4) or upgradeability mechanism (Section 4.5), ensuring correct operation of the forced queue. Moreover, the described upgradeability mechanism also ensures that all users have a fair chance to process their transactions before new policies take effect. Notably, if we allow the query that updates the blacklisting policy “on-the-spot”, then Alloy immediately generates simple counterexamples to most of the properties mentioned above. Most importantly, due to the non-deterministic order of the actions in the system, “on-the-spot” updates can always lead to a situation where forced queue inputs get blacklisted, which makes the system stuck (i.e., no more blocks could be finalized).
@@
Potential vulnerability in state-of-the-art upgradeability policy. The current state-of-the-art, often described as the safest approach, typically includes one week before enforcing an upgrade. [^13] https://medium.com/l2beat/introducing-stages-a-framework-to-evaluate-rollups-maturity-d290bb22befe Upgrades based solely on time can lead to situations where users who wish to exit the L2 before the upgrade are locked in because their transactions cannot be processed promptly. However, our solution (Section 4.5) mitigates this issue by ensuring that all transactions in the forced queue are processed before the upgrade takes effect, thus maintaining users’ safety in case of malicious upgrades. Here, we discuss the potential pitfalls of relying solely on a timeout period for upgrades and then provide a counterexample demonstrating the issues with that mechanism, as produced by the Alloy Analyzer.

### v3

@@ 3.1. Threat Model
Users. Users interact with the L2 either by submitting their transactions directly to the L2 network, or by submitting them through the rollup’s contract on the L1 via the forced queue.
@@
Rollup administrators. In addition to rollup operators, some entities have administrator capabilities for the L1 smart contracts of the L2. Administrators can be white-listed or elected via governance mechanisms. These administrators have the ability to request system upgrades or update blacklists. In practice, the administrator capabilities are assigned to a multisig that could be controlled by some distinct entities. The mechanisms we describe in this paper should maintain security even if administrators are malicious.
@@
  • The cryptographic primitives (ZKPs) and their implementation are assumed to be secure (Chaliasos et al., 2024b), ensuring that no verifier accepts an invalid proof of execution. This assumption also extends to the correctness of any trusted setup procedures, the deployed verification keys on the L1 verifier contract, and the smart contract code responsible for proof verification.
  • Smart contracts operating L2 functionalities on the L1 are considered secure against conventional vulnerabilities, such as re-entrancy attacks (Chaliasos et al., 2024a), and are modifiable only through a well-defined upgrade process. In other words, in this paper, we focus on what may be called logical vulnerabilities.
@@
Strong Adversary Model. Any of the parties mentioned above — users, operators, and administrators — can become malicious. Users can perform DoS attacks, operators can engage in transaction censorship, and rollup administrators can execute malicious upgrades that could lead to the theft of all user assets. Despite these possibilities, the saving grace is that rollup contracts deployed on the underlying L1, together with the economic security that the L1 provides, allow us to use decentralized trust to address these issues. This comes down to the ability to analyze (Firsov and Livshits, 2024) these rollup contracts.
For example, zkSync Era has contracts such as Verifier.sol, which process validity proofs and can be upgraded by a multisig [^12] https://github.com/matter-labs/era-contracts. It is critical that this verifier can only be updated through a safe process because if it becomes malicious, users can lose their money. Hence, users should have a sufficient mechanism to exit the rollup in case of adversarial or questionable upgrades. While our model is extensible, this paper largely focuses on functional security, sidestepping DoS concerns.
3.2. User-centric Security Goals
@@
Weak liveness. A key goal is to ensure that any user transaction submitted to the forced queue is processed as long as the L2 continues to make progress. This guarantee is crucial to protecting users against potentially compromised or malicious L2s or even adding L1-like censorship resistance for L2s. Moreover, our proposed design ensures that even when L2 contracts enforce transaction blacklisting, transactions cannot be halted if they have already been submitted to the forced queue. This mechanism is vital to maintaining the integrity and safety of ZK-Rollups. In summary, weak liveness means that the L2 will eventually process user transactions submitted through L1 (i.e., inheriting the liveness of the L1), or the rollup will be frozen. In Section 4.3, we design a forced queue mechanism that enforces weak liveness.
@@
4. Implementation in Alloy
In this section, we present the implementation details of our ZK Rollup model using the Alloy specification language. We first give a brief technical introduction to Alloy, followed by the Alloy model for a strawman ZK Rollup and three critical mechanisms, highlighting our reasoning and essential security properties. In Section 5.1, we demonstrate how to use our model to detect issues in flawed designs. Our approach directly encourages finding problems early, at the design stage, as opposed to later.
Why Alloy. We selected Alloy 6 (Brunel et al., 2021) over other formal methods tools, such as TLA+ and NuSMV, due to its balance of expressiveness, ease of specification, and automation for finding counterexamples. Alloy’s declarative syntax based on first-order relational logic is particularly well-suited for modeling relational systems, such as rollup states and transitions (Jackson, 2002). Prior work in web security (Akhawe et al., 2010) and smart contract validation (Godoy et al., 2022) shows that Alloy enables uncovering subtle vulnerabilities effectively, making it a practical and proven choice for the security analysis of systems where quick design iteration and counterexample-driven debugging are critical.
4.1. A Brief Introduction to Alloy
Alloy 6 (Jackson, 2002; Brunel et al., 2021) is a declarative modeling language that uses first-order relational logic to specify and check the properties of software systems. Data types in Alloy are represented as relations defined by signatures, which can be abstract or concrete. Abstract signatures categorize general types that are refined by more specific concrete signatures, creating a flexible and powerful type system. The language enforces the integrity of the model through facts, which are constraints that must always hold within the model. Functions and predicates in Alloy are named expressions and logical formulas, respectively; these allow the parametrization and reuse of complex logical constructs throughout the model.
@@ 4.1. A Brief Introduction to Alloy
The Alloy analysis is always limited to a specified scope. The small-scope hypothesis states that a high proportion of errors can be found by analyzing a system for all test inputs within some small scope (Jackson and Damon, 1996; Jackson, 2019). However, Alloy’s guarantees are strictly bounded: if no counter-example appears within the specified scope, nothing can be inferred beyond it. To offset this incompleteness, we rely on the small-scope hypothesis. Still, bounded success is not a universal proof. Alloy is a performant bug-finder, not a full verifier. We discuss this trade-off further in Section 6.
This structured approach to modeling with Alloy ensures model checking of properties and behaviors in software systems, making it an invaluable tool in the development and analysis of reliable software architectures. In the upcoming sections, we present how we used Alloy to design and verify the correctness of critical elements of ZK-Rollup logic.
4.2. Strawman ZK Rollup Model
@@
    3 c.state.subseq[0,sub[#L1.finalized_state,1]] == L1.finalized_state
    4 L1.commitments == L1.commitments + c
@@
    4 c.state == p.state
    5 c.diff == p.diff
    6 c.state == L1.finalized_state
@@
    8 L1.finalized_state == L1.finalized_state.add[p.diff]
@@
7 c.diff == p.diff and b == c.diff
8 and c.state == L1.finalized_state)
@@
Strictly speaking when we checked the property for 20 steps it still can fail for 21 steps. However, small-scope hypothesis states that a high proportion of errors can be found by checking a program for all test inputs within some small scope (Jackson and Damon, 1996).
4.3. Forced Queue
@@
3 L1.forced_queue == L1.forced_queue.add[f]
@@
5 and not L1.forced_queue.first.tx == L1.forced_queue’.first.tx
@@
4.4. Forced Queue with Blacklisting
Recently, transaction and address censorship have become prominent issues in the Ethereum ecosystem. For example, 6% of Ethereum blocks are constructed to be OFAC-compliant (Wahrstätter et al., 2024), and Circle (USDC) maintains a list with blacklisted addresses (Wang et al., 2023). Some L2 networks may choose to censor transactions to be OFAC compliant, potentially to be determined through their governance protocol. Next, we describe an “enshrined” blacklisting mechanism; this mechanism was meticulously designed, formalized, and checked with Alloy Analyzer to ensure it does not violate forced queue properties.
@@ 4.4. Forced Queue with Blacklisting
2 L1.forced_queue.first == f
3 L1.blacklist == L1.forced_queue.first.predicate
4 L1.forced_queue == L1.forced_queue.delete[0]
@@
4.5. Forced Queue and Upgradeability
@@
  4. (4)
    Once the forced queue is cleared, the upgrade is deployed, and the rollup shifts to a new operational mode. During this timeout period, only forced queue events are processed, which pressures the L2 operator to handle the forced queue swiftly. Halting the acceptance of forced transactions risks the assurance of a permissionless exit via the forced queue mechanism. Note that L2 users can still submit transactions through the L2 and receive pre-confirmations, but these cannot be finalized on the L1 until the forced queue is cleared.
@@
In summary, both blacklisting strategies have their trade-offs. More specifically, the blacklisting strategy through the forced queue (see Section 4.4) allows the ZK-Rollup to immediately prevent censored transactions from being added into the forced queue (however, it cannot forbid execution of the transactions that are already present in the queue). On the other hand blacklisting strategy through the upgradeability mechanism allows users to notice the upgrade and quit the ZK-Rollup within the predefined window of time without receiving a damage.
@@
2 L1.ongoing_upgrade == none
3 L1.ongoing_upgrade == f
@@
3 not L1.blacklist == L1.blacklist implies
@@
5 once (L1.ongoing_upgrade == none and
6 L1.ongoing_upgrade == x and
7 L1.blacklist == is and
@@
2 (not L1.blacklist == L1.blacklist’) implies
@@
Given all of the above properties, we arrive at the following:
@@
This mechanism ensures that users are not caught off-guard by upgrades and are guaranteed to be able to act upon upgrade (e.g., exit from L2 rollup) before the upgrade is enforced.
Figure 3. Counter-example generated by Alloy, illustrating potential issues with the naive upgradeability mechanism. In this scenario, the blacklisting policy targets Input_1, leading to an L2 freeze because an input in the forced_queue cannot be processed. This occurs because the mechanism does not require processing all transactions from the forced queue prior to executing an upgrade.
5. Leveraging Alloy Models for Testing Rollups
This section presents a complete methodology for testing and verifying the correctness of ZK-Rollup designs and implementations. First, in Section 5.1, we describe how our Alloy models can be used for design-time testing, enabling early detection of security and censorship resistance issues in the system design. Then, in Section 5.2, we explain how the formal properties we specify in Section 4 (i.e., FQPs, BPs, and UPs) can be adapted for code-level verification and testing of smart contracts, using industry-standard formal verification and property-based testing frameworks. Further, we demonstrate this methodology for Scroll’s forced queue, where we adapt our Alloy model to reflect Scroll’s architecture and translate the formal properties into Foundry invariants. This allows us to apply property-based testing directly on Scroll’s smart contracts, bridging design-time specifications with implementation-level testing. Together, these two layers of testing and verification form a practical end-to-end methodology for testing the correctness of ZK-Rollup security mechanisms from the initial design to the deployed code.
Note on state-of-the-art rollups. Currently, most general-purpose ZK-Rollups lack implementations of essential mechanisms like forced queues and delayed upgrades, which are crucial for inheriting L1’s censorship resistance and security. The only exception is Scroll that implements a timeout-based forced queue that we thoroughly analyze in this section. Yet, our approach can guide practitioners in testing their implementations as these mechanisms are developed.
5.1. Design-time Testing
Building on the implementation details and security properties described in Section 4, we explore how our Alloy models can be used for design-time testing of ZK-Rollups. We provide an infeasibility result and highlight a vulnerability in the design of upgradeability mechanisms for ZK-Rollups. Note that this is not affecting any live systems as the safe upgraedability functionality has not be deployed by any ZK-Rollups.
@@ 5.1. Design-time Testing
In this scenario for the rollup with upgradeability and blacklisting (Section 4.5), we start with an empty blacklist state and no queued commitments or proofs. We assert that, eventually, we would like to reach a state where there is an input x which is blacklisted but not finalized. After that we would like to eventually reach a state when the input x is finalized. This scenario is only possible if the rollup updates its blacklists two times where the first update blacklists x and the second update removes x from the blacklist and at the end x gets finalized. When we execute the run command, the Alloy Analyzer looks for models with up to seven instances of each signature that would satisfy the predicate. The Alloy Analyzer finds a satisfying instance that consists of a trace of length eight.
@@
Infeasibility of on-the-spot blacklisting. When processing a blacklisting policy update transaction (Section 4.4), transactions that precede it in the queue must be processed. This means that otherwise censored transactions would have to be processed. If blacklisting is applied immediately, it can generate a trace that gets the system stuck, breaking the guarantees of our system. To maintain the integrity of the forced queue, blacklisting should be enforced either through the forced queue (Section 4.4) or upgradeability mechanism (Section 4.5), ensuring correct operation of the forced queue. Moreover, the described upgradeability mechanism also ensures that all users have a fair chance to process their transactions before new policies take effect. Notably, if we allow the query that updates the blacklisting policy “on-the-spot”, then Alloy immediately generates simple counterexamples to most of the properties mentioned above. Most importantly, due to the non-deterministic order of the actions in the system, “on-the-spot” updates can always lead to a situation where forced queue inputs get blacklisted, which makes the system stuck (i.e., no more blocks could be finalized).
@@
Potential vulnerability in upgradeability policy. The current state-of-the-art upgradeability design, often described as the safest approach, typically includes one week before enforcing an upgrade. [^14] https://medium.com/l2beat/introducing-stages-a-framework-to-evaluate-rollups-maturity-d290bb22befe Upgrades based solely on time can lead to situations where users who wish to exit the L2 before the upgrade are locked in because their transactions cannot be processed promptly. However, our solution (Section 4.5) mitigates this issue by ensuring that all transactions in the forced queue are processed before the upgrade takes effect, thus maintaining users’ safety in case of malicious upgrades. Here, we discuss the potential pitfalls of relying solely on a timeout period for upgrades and then provide a counterexample demonstrating the issues with that mechanism, as produced by the Alloy Analyzer.